OpenStudy (anonymous):

CAN SOMEONE HELP ME PLEASE! I WILL FAN YOU AND GIVE YOU A MEDAL ON EVERYTHING!!!! what are the factored forms of the following expressions? 1. 6x^2+5x+1 2.6x^2+17x+12 3.12d^2+4d-1 4.8g^2+6g-9 5.20x^2+22x-12

OpenStudy (anonymous):

u can use this formula x=-b+/- sqrt (b^2-4ac) divided 2a...or u can use calculator..

OpenStudy (anonymous):

i just don't understand it. can you please help me?

OpenStudy (anonymous):

\[x=(-b \pm \sqrt{b ^2-4ac})/2a\]

OpenStudy (anonymous):

i cant enter that in a calculator.

OpenStudy (anonymous):

if using calculator i get x=-1/3 or x=-1/2 for question number 1 ....

OpenStudy (anonymous):

that isn't an answer choice tho. if i give you the answer choices will that help?

OpenStudy (anonymous):

if i can help u....i will....can give me the answer...

OpenStudy (anonymous):

for 1 A.(3x-1)(2x-1) B.(3x+1)(2x-1) C.(3x-1)(2x+1) D.(3x+1)(2x+1)

OpenStudy (anonymous):

owh...like that....the answer..just same what i do it....i get x=-1/3 and x=-1/2...right....wait...i will proof that to u....and if u dont understands...just tell me..ok? ^_^...

OpenStudy (anonymous):

for 2 A.(3x-4)(2x+3) B.(3x-4)(2x-3) C.(3x+4)(2x+3) D.(3x+4)(2x-3)

OpenStudy (anonymous):

so for 1 what would the answer be?

OpenStudy (anonymous):

but i get (3x+1)(2X+1)...... all plus....

OpenStudy (anonymous):

so its D?

OpenStudy (anonymous):

if you help me i will give you as many medals as i can

OpenStudy (anonymous):

\[6x^2+5x+1=0\]

OpenStudy (anonymous):

\[(3x+1)(2x+1)=0\]

OpenStudy (anonymous):

so is D right or wrong?

OpenStudy (anonymous):

right...the answer is D

OpenStudy (anonymous):

ok what about number 2

OpenStudy (anonymous):

C

OpenStudy (anonymous):

for 4 A.(4g+3)(2g+3) B.(4g-3)(2g-3) C.(4g+3)(2g-3) D.(4g-3)(2g+3)

OpenStudy (anonymous):

\[3. (6x+1)(6x+1) or (6x+1)^2\]

OpenStudy (anonymous):

for 5 A.2(5x-2)(2x+3) B.2(5x+2)(2x-3) C.(10x-2)(4x+3) D.2(5x+4)(2x-3)

OpenStudy (anonymous):

4. D

OpenStudy (anonymous):

what about 5 and then i swear i will give you many medals

OpenStudy (anonymous):

5. A....

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

i made a 100! thank you!

OpenStudy (anonymous):

welcome...